OpenStudy (anonymous):

What is the solution to the equation? 2 - z/3 = 12 A. z = –42 B. z = –34 C. z = –30 D. z = 42

OpenStudy (anonymous):

start by subtract ing 2 from both sides second step, multiply both sides by -3

OpenStudy (anonymous):

I got -36

OpenStudy (anonymous):

Is the answer B?

OpenStudy (anonymous):

\[2-\frac{z}{3}=12\] Subtract 2 from both sides\[2-\frac{z}{3}-2=12-2\] The twos cancel on the left\[-\frac{z}{3}=12-2\] Multiply both sides by -3\[-3\left( -\frac{z}{3} \right)=-3(12-2)\] The negative threes cancel on the left \[z=-3(12-2)\] subtract 2 from 12, then multiply by -3

OpenStudy (anonymous):

So the answer is C.

OpenStudy (anonymous):

Yep.